OpenStudy (colouringme):

how to proof that 5^x times 4 = 25 times 5^x I think of logarithm but I don't know

ganeshie8 (ganeshie8):

\(\large 5^x \times 4 = 25\) or \(\large 5^{x \times 4} = 25\)

ganeshie8 (ganeshie8):

which one is it /

ganeshie8 (ganeshie8):

?

OpenStudy (colouringme):

the equation won't show up in my laptop.

ganeshie8 (ganeshie8):

http://prntscr.com/34izzp

OpenStudy (colouringme):

\[5^{x} \times 4 + 1 = 5^{x} \times 5\] sorry, it suppose to be this.

ganeshie8 (ganeshie8):

easy, then

ganeshie8 (ganeshie8):

move 5^x * 4 to right hand side : 1 = 5^x*5 - 5^x*4

OpenStudy (colouringme):

my gosh, why I didn't think about it. THANK YOU! :D

ganeshie8 (ganeshie8):

that gives : 5^0 = 5^x or x = 0 is the solution for 5^x+1 = 5^x*5

ganeshie8 (ganeshie8):

u wlc :)